# Robust variational neural posterior estimation for simulation-based inference

### v1

@@ Abstract
Recent advances in neural density estimation have enabled powerful simulation-based inference (SBI) methods that can flexibly approximate Bayesian inference for intractable stochastic models. Although these methods have demonstrated reliable posterior estimation when the simulator accurately represents the underlying data generative process (GDP), recent work has shown that they perform poorly in the presence of model misspecification. This poses a significant problem for their use on real-world problems, due to simulators always misrepresenting the true DGP to a certain degree. In this paper, we introduce robust variational neural posterior estimation (RVNP), a method which addresses the problem of misspecification in amortised SBI by bridging the simulation-to-reality gap using variational inference and error modelling. We test RVNP on multiple benchmark tasks, including using real data from astronomy, and show that it can recover robust posterior inference in a data-driven manner without adopting tunable hyperparameters or priors governing the misspecification.
@@ 1 Introduction
Simulator models are ubiquitous in many areas of the natural sciences and engineering, enabling researchers to approximate complex real-world data-generating processes (DGP) using physically grounded forward models. However, these simulators are often computationally expensive, non-differentiable, and lack closed-form likelihoods, making traditional inference methods inapplicable. Implicitly, the simulator defines an intractable likelihood $p(\bm{x}_{\rm sim}|\,\bm{\theta})$ over $\mathcal{X}_{\rm sim}\subseteq\mathbb{R}^{n}$, relating the simulated observations and the parameters of interest $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{m}$. As a result of intractability and the computational expense of running simulations, solving the inverse problem of inferring simulator parameters from observed data $\bm{x}_{\rm obs}$ poses a significant challenge. Simulation-based inference (SBI, Cranmer et al. 2020) or likelihood-free inference provides methods to approximately infer the posterior distribution of the simulator parameters conditioned on observed data.
@@
A range of SBI methods have emerged to solve the likelihood-free inference problem, beginning with traditional approaches such as approximate Bayesian computation (ABC; Rubin 1984; Beaumont et al. 2002) and Bayesian synthetic likelihood (BSL, Price et al. 2018). Recent work has introduced methods based on neural density estimation, such as neural posterior estimation (NPE, Papamakarios & Murray 2016; Lueckmann et al. 2017; Greenberg et al. 2019), neural likelihood estimation, (NLE, Lueckmann et al. 2019; Papamakarios et al. 2019), neural ratio estimation (NRE, Izbicki et al. 2014; Cranmer et al. 2016; Hermans et al. 2020; Durkan et al. 2020), and diffusion-based methods (Glöckler et al., 2024). SBI methods can be categorised into amortised and non-amortised inference methods. In the context of neural SBI, non-amortised methods such as sequential neural posterior estimation (SNPE; Greenberg et al. 2019), sequential neural likelihood estimation (SNLE; Papamakarios et al. 2019), and sequential neural ratio estimation (SNRE; Hermans et al. 2020) target a single posterior conditioned on fixed data, adapting their inference procedure with each simulation round. After an up-front simulation budget, amortised methods aim to learn a global inference model over a given prior, making them well-suited for scenarios where repeated or scalable inference is required. In this paper, we focus on amortised SBI methods for inferring posterior distributions for any observation within the support of the simulator model.
@@
SBI methods have been widely used in fields such as astronomy (Mishra-Sharma & Cranmer, 2022), particle physics (The Atlas Collaboration, 2025), cosmology (Lemos et al. 2023; Zeghal et al. 2024), and neuroscience (Oesterle et al. 2020; Hashemi et al. 2024), to name a few. However, recent work has shown that they can yield overconfident posterior approximations (Hermans et al., 2022) and suffer significantly when the true DGP does not lie within the family of distributions defined by the statistical model (Cannon et al. 2022; Schmitt et al. 2024), known as model misspecification. Model misspecification may be caused by a variety of factors, such as contamination in the data or unaccounted-for physical processes in the modelling that can lead to overconfident posteriors (Hermans et al., 2022). This discrepancy between the simulated data and the real observations is known as the simulation-to-reality gap (Miglino et al., 1995) or simulation gap.
@@
Methods for mitigating against misspecification in neural SBI have done so mainly by addressing the simulation-to-reality gap. This is based on the assumption that misspecification appears as a divergence-based discrepancy between the true DGP $p^{*}(\bm{x}_{\rm obs})$ and the distribution described by the simulator model $p(\bm{x}_{\rm sim})=\mathbb{E}_{p(\bm{\theta})}[p(\bm{x}_{\rm sim}|\,\bm{\theta})]$, under the prior distribution $p(\bm{\theta})$. Robust SBI methods usually address the simulation-to-reality gap through error modelling and adjustment parameters (Ward et al. 2022; Frazier & Drovandi 2021; Kelly et al. 2024), domain adaptation approaches (Huang et al. 2023; Swierc et al. 2024; Elsemüller et al. 2025; Mishra et al. 2025), or generalized Bayesian inference (Dellaporta et al., 2022). The success of most of these methods relies on the observed points appearing as out-of-distribution (OOD) with respect to the simulated observations. However, recent work has underscored the importance of within-distribution (ID) points in a misspecified SBI (Schmitt et al. 2024;Frazier et al. 2024;Elsemüller et al. 2025) as the errors in the model may still produce summary statistics which lie ID relative to the simulations. Wehenkel et al. (2025) shows that using a reliable calibration set can aid towards robust amortised SBI under such modelling errors. Often, a reliable calibration set will not exist, making such problems highly difficult to solve. Recently, unsupervised domain adaptation (UDA) methods have been implemented in robust amortised SBI using Maximum Mean Discrepancy and domain-adversarial neural networks (Elsemüller et al., 2025), and using consistency loss regularisation (Mishra et al., 2025). As amortised SBI looks to construct general posteriors for a range of observations, it is natural to consider the misspecification problem for situations involving many observations where all points appear OOD, or when a significant number of points appear OOD.
@@
Despite their success in robust SBI, existing methods for robust SBI encounter issues in the context of robust amortised SBI. In particular, the error modelling and correction parameter approaches scale poorly to amortised Bayesian inference due to their dependence on an MCMC sampling step. On the other hand, they benefit from their Bayesian formulation, particularly through the connection between hyperparameter choice and Bayesian prior adoption (Ward et al. 2022; Frazier & Drovandi 2021; Kelly et al. 2024). Domain adaptation methods scale more favourably to amortised SBI, but come at the cost of a non-Bayesian interpretation of the domain adaptation hyperparameters, a lack of interoperability of the domain adaptation (Elsemüller et al., 2025), and a lack of clarity between the trade-off in the domain adaptation and the inference algorithm (Chen et al., 2021). Furthermore, it is not always desirable to use domain-adapted neural embedding statistics if expert knowledge on the summary embedding space is available, such as known sufficient statistics on a low-dimensional observation space in physically motivated units. Data-driven methods that have a reliable Bayesian interpretation and do not rely on hyperparameters are desirable in this context.
@@ 1.1 Our contributions
We propose robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T) that addresses the misspecification problem in amortized SBI by pre-training the simulator likelihood $p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$ for parameters $\bm{\Psi}$, adopting a flexible error model $p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})$, and using an importance weighted autoencoder (Burda et al., 2015) scheme to maximize the evidence of the true data under the variational posterior $p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ for the parameters $\bm{\xi}$ and $\bm{\phi}$. The variational posterior is the main objective of the RVNP algorithm. The posterior $p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ is optionally tuned on the adapted synthetic DGP to return the RVNP-T posterior distribution. The error modelling we choose to adopt can be seen as a neural adaptation of the mean adjustment and covariance inflation from Frazier & Drovandi (2021).
The main claim of our paper is that robust variational neural posterior estimation (RVNP) and its tuned variant can recover robust amortised posterior inference under misspecification by bridging the simulation-to-reality gap using error modelling. The error model parameters are adapted in a data-driven way when we have many observations. We summarise our contributions as follows:
@@
    We introduce RVNP, an amortised SBI method that uses a pre-trained simulator likelihood, an error model, and an importance-weighted autoencoder (Burda et al., 2015) scheme to return robust amortised posterior inference under misspecification without adopting tunable parameters or priors over the misspecification. We also introduce RVNP-T, which tunes the final posterior using the simulator and the noise induced by the error model.
@@
    To our knowledge, this is the first example of using amortised variational autoencoders to address the misspecification problem in SBI.
In amortised SBI, it is usually necessary to have a pre-defined up-front simulator sample size, which may change if the model is learnt in rounds. We assume that we have a fixed simulation budget and that the neural statistic embedding is either pre-trained or adopted by expert knowledge. The simulation-to-reality gap should inform our error model as we observed more data from the true DGP (Figure 1). We impose a strong inductive bias on the domain adaptation so that the error model can only inflate the synthetic DGP to account for the simulation gap, allowing us to have a form of model criticism. Cranmer et al. (2020) suggests augmenting the simulator with an error model to account for the simulation-to-reality gap. However, choosing an error model can be difficult and arbitrary and requires multiple independent trainings to test the different error models (Ward et al., 2022). We instead consider using an error model that is inferred by maximising the evidence lower bound jointly with inferring the posterior distribution of the parameters. Our method is well-suited to situations when the misspecified points appear, on average, OOD relative to the simulated points. Current unsupervised robust SBI methods rely on a clear simulation-to-reality gap driven by misspecification. In situations where the entire inference set lies within or very near the original simulated distribution, a strong calibration set is most likely necessary.
@@ 2.1 SBI formalism
We let $\iota_{\omega}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{l}$, $\bm{x}_{\rm sim}\mapsto\bm{z}_{\rm sim}$ denote a statistical embedding parametrized by $\omega$. This embedding can represent fixed user-defined summary statistics, a pre-defined embedding, or a neural statistic estimator (NSE) where the parameters $\bm{\omega}$ are to be learnt. Lower-dimensional embeddings are important when dealing with high-dimensional data, but can come at the cost of information loss when the embedding is not a sufficient statistic of $\bm{x}_{\rm sim}$ for $\bm{\theta}$ (Blum et al., 2013).
@@ 2.2 Amortised neural posterior and neural likelihood estimation
The goal of amortised neural posterior estimation is to approximate the unknown posterior distribution $p^{*}(\bm{\theta}|\bm{x_{\rm obs}})$ for all $\bm{x}_{\rm obs}\in\mathcal{X}_{\rm obs}$. After choosing a conditional density estimation architecture $p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ parametrized by $\bm{\phi}$, and an architecture for the neural statistic embedding, $\iota_{\omega}$, NPE (Papamakarios & Murray 2016; Lueckmann et al. 2017; Greenberg et al. 2019) fits for the parameters $\bm{\omega}$ and $\bm{\phi}$ by minimizing the the expected forward Kullback-Leibler (KL) divergence between analytic and approximate posterior
$\mathcal{L}_{\rm NPE}({\bm{\phi},\bm{\omega}})$ $=\mathbb{E}_{p^{*}(x_{\rm obs})}\big{[}\mathbb{KL}[p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))||p_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big{]}$ (1)
$=\mathbb{E}_{p^{*}(x_{\rm obs})}\big{[}\mathbb{E}_{p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))}[-\log p_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big{]},$
where the expectation is over the unknown true data-generating distribution $p^{*}(\bm{x}_{\rm obs})$ and the second line follows because the true unknown posterior and its entropy do not depend on the trainable parameters. As noted in Schmitt et al. (2024), this amortised posterior objective function is not feasible as we may not have enough real data to approximate the expectation with respect to $p^{*}(\bm{x}_{\rm obs})$, and the true posterior is unknown and intractable. Instead, the unknown $p^{*}(\bm{x}_{\rm obs})$ is replaced with the marginal likelihood $p(\bm{x_{\rm obs}})=\int p(\bm{x}_{\rm obs}|\,\bm{\theta})p(\bm{\theta})d\bm{\theta}$. Under the model assumption we have $\bm{x}_{\rm obs}$ can be replaced with $\bm{x}_{\rm sim}$ and the objective becomes
$\mathcal{L}(\bm{\phi},\bm{\omega})_{\rm NPE}:=-\mathbb{E}_{\bm{x}_{\rm sim},\bm{\theta}}[\log p_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm sim}))].$ (2)
This is minimised with respect to the parameters $\bm{\omega}$ and $\bm{\phi}$. The success of this objective depends on the assumption that sampling from the evidence is equivalent to sampling from $p^{*}(\bm{x}_{\rm obs})$.
On the other hand, neural likelihood estimation (NLE) learns a distribution maximising the conditional log-probability of the simulated data
@@
with respect to the flow parameters $\bm{\Psi}$ (where we dropped reference to $\iota_{\bm{\omega}}$), which is equivalent to minimising the KL divergence between the flow and the target distribution. The assumption of the learnt distribution being a likelihood identifies $\bm{x}_{\rm obs}$ with $\bm{x}_{\rm sim}$. However, one can also view the NLE objective as a surrogate training objective to approximate the sampling distribution given by the simulator, without any initial assumptions about the connection between true data and simulator output. In this sense, the NLE objective becomes a neural conditional prior proxy, and a likelihood can be introduced to relate the forward model to the true $\bm{x}_{\rm obs}$.

### v2

# Misspecification-robust amortised simulation-based inference using variational methods

@@ Abstract
Recent advances in neural density estimation have enabled powerful simulation-based inference (SBI) methods that can flexibly approximate Bayesian inference for intractable stochastic models. Although these methods have demonstrated reliable posterior estimation when the simulator accurately represents the underlying data generative process (DGP), recent work has shown that they perform poorly in the presence of model misspecification. This poses a significant issue for their use in real-world problems, due to simulators always misrepresenting the true DGP to a certain degree. In this paper, we introduce robust variational neural posterior estimation (RVNP), a method which addresses the problem of misspecification in amortised SBI by bridging the simulation-to-reality gap using variational inference and error modelling. We test RVNP on multiple benchmark tasks, including using real data from astronomy, and show that it can recover robust posterior inference in a data-driven manner without adopting hyperparameters or priors governing the misspecification influence.
@@ 1 Introduction
Simulator models are ubiquitous in many areas of the natural sciences and engineering, enabling researchers to approximate complex real-world data generative processes (DGP) using physically grounded forward models. However, these simulators are often computationally expensive, non-differentiable, and lack closed-form likelihoods, making traditional inference methods inapplicable. Implicitly, the simulator defines an intractable likelihood $p(\bm{x}_{\rm sim}|\,\bm{\theta})$ over $\mathcal{X}_{\rm sim}\subseteq\mathbb{R}^{n}$, relating the simulated observations and the parameters of interest $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{m}$. As a result of intractability and the computational expense of running simulations, solving the inverse problem of inferring simulator parameters from observed data $\bm{x}_{\rm obs}$ poses a significant challenge. Simulation-based inference (SBI, Cranmer_2020) or likelihood-free inference provides methods to approximately infer the posterior distribution of the simulator parameters conditioned on observed data.
@@
A range of SBI methods have emerged to solve the likelihood-free inference problem, beginning with traditional approaches such as approximate Bayesian computation (ABC; 10.1214/aos/1176346785; abc_genetics) and Bayesian synthetic likelihood (BSL, Price02012018). Recent work has introduced methods based on neural density estimation, such as neural posterior estimation (NPE, 2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree), neural likelihood estimation (NLE, pmlr-v96-lueckmann19a; papamakarios2019sequentialneurallikelihoodfast), neural ratio estimation (NRE, pmlr-v33-izbicki14; cranmer2016approximatinglikelihoodratioscalibrated; hermans2020likelihoodfreemcmcamortisedapproximate; pmlr-v119-durkan20a), and modelling the joint distribution of data and simulation parameters (DBLP:conf/icml/GlocklerDWWM24). SBI methods can be categorised into amortised and non-amortised inference methods. In the context of neural SBI, non-amortised methods such as sequential neural posterior estimation (SNPE; 2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree), sequential neural likelihood estimation (SNLE; papamakarios2019sequentialneurallikelihoodfast), and sequential neural ratio estimation (SNRE; hermans2020likelihoodfreemcmcamortisedapproximate) target a single posterior conditioned on fixed data, adapting their inference procedure with each simulation round. After an up-front simulation budget, amortised methods aim to learn a global inference model over a given prior, making them well-suited for scenarios where repeated or scalable inference is required. In this paper, we focus on SBI methods for inferring posterior distributions that are amortised over a dataset of observations.
@@
SBI methods have been widely used in fields such as astronomy (PhysRevD.105.063017), particle physics (atlascollaboration2024implementationneuralsimulationbasedinference), cosmology (Lemos_2023; zeghal2024simulationbasedinferencebenchmarklsst), and neuroscience (10.7554/eLife.54997; 2024MLS&T...5c5019H), to name a few. However, recent work has shown that they can yield overconfident posterior approximations (2021arXiv211006581H) and suffer significantly when the true DGP does not lie within the family of distributions defined by the statistical model (cannon2022investigating; schmitt2024detecting), known as model misspecification. Model misspecification may be caused by a variety of factors, such as contamination in the data or unaccounted-for physical processes in the modelling that can lead to overconfident posteriors (2021arXiv211006581H). This discrepancy between the simulated data and the real observations is known as the simulation-to-reality gap (miglino1995evolving) or simulation gap.
@@
Methods for mitigating against misspecification in neural SBI have done so mainly by addressing the simulation-to-reality gap. This is based on the assumption that misspecification appears as a divergence-based discrepancy between the true DGP $p^{*}(\bm{x}_{\rm obs})$ and the marginal distribution described by the simulator model $p(\bm{x}_{\rm sim})=\mathbb{E}_{p(\bm{\theta})}[p(\bm{x}_{\rm sim}|\,\bm{\theta})]$, under the prior distribution $p(\bm{\theta})$. Robust SBI methods often address the simulation-to-reality gap through error modelling and adjustment parameters (ward2022robust; frazier2021robust; kelly2024misspecificationrobust), domain adaptation approaches (huang2023learning; swierc2024domain; elsemüller2025doesunsuperviseddomainadaptation; mishra2025robust), or generalised Bayesian inference (dellaporta2022robust). The success of most of these methods relies on the observed points appearing as out-of-distribution (OOD) with respect to the simulated observations. However, recent work has underscored the importance of within-distribution (ID) points in a misspecified SBI (schmitt2024detecting; Frazier26092024; elsemüller2025doesunsuperviseddomainadaptation), as the errors in the model may still produce summary statistics which lie ID relative to the simulations. wehenkel2025addressing showed that using a reliable calibration set can aid towards robust amortised SBI under such modelling errors. Often, a reliable calibration set will not exist, making such problems highly difficult to solve. Recently, unsupervised domain adaptation (UDA) methods have been implemented in robust amortised SBI using maximum mean discrepancy (MMD), domain-adversarial neural networks (elsemüller2025doesunsuperviseddomainadaptation), and consistency loss regularisation (mishra2025robust). As amortised SBI looks to construct general posteriors for a range of observations, it is natural to consider the misspecification problem for situations involving many observations where all points appear OOD, or when a significant number of points appear OOD.
@@
Despite their success in robust SBI, existing methods encounter issues in the context of robust amortised SBI. In particular, the error modelling and correction parameter approaches scale poorly to amortised Bayesian inference due to their dependence on a Markov Chain Monte Carlo (MCMC) sampling step. On the other hand, they benefit from their Bayesian formulation, particularly through the connection between hyperparameter choice and Bayesian prior adoption (ward2022robust; frazier2021robust; kelly2024misspecificationrobust). Domain adaptation methods scale more favourably to amortised SBI, but come at the cost of a non-Bayesian interpretation of the domain adaptation hyperparameters, a lack of interoperability of the domain adaptation (elsemüller2025doesunsuperviseddomainadaptation), and a lack of clarity between the trade-off in the domain adaptation and the inference algorithm (chen2021neuralapproximatesufficientstatistics). Furthermore, it is not always desirable to use domain-adapted neural embedding statistics if expert knowledge on the summary embedding space is available, such as known sufficient statistics on a low-dimensional observation space in physically motivated units. Data-driven methods that have a reliable Bayesian interpretation and do not rely on hyperparameters are desirable in this context.
@@ 1.1 Our contributions
Problem setting: Given a dataset $\mathcal{O}=\{\bm{x}_{\rm obs}^{(i)}\}_{i=1}^{N_{\rm obs}}$ where each observation corresponds to a different unknown parameter $\bm{\theta}^{*(i)}$, we seek to infer all posterior distributions $\{p(\bm{\theta}\,|\,\bm{x}_{\rm obs}^{(i)})\}_{i=1}^{N_{\rm obs}}$, when the simulator model is misspecified.
Amortised SBI: In this paper, we refer to amortised SBI as amortising the Bayesian inference over all data points in $\mathcal{O}$. This definition is in line with amortised variational inference (margossian2023amortised).
Our solution: We propose robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T), methods that address the misspecification problem in amortised SBI. RVNP pre-trains a simulator likelihood $p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$, adopts a flexible error model $p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})$, and uses an importance weighted autoencoder (burda2015importance) inference scheme to maximize the evidence of the true data under the variational posterior $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ for the parameters $\bm{\xi}$ and $\bm{\phi}$.
The main claim of our paper is that RVNP and its tuned variant can recover robust amortised posterior inference under misspecification by bridging the simulation-to-reality gap using error modelling. The error model parameters are adapted in a data-driven way when we have many observations and the method requires no extra manually tuned misspecification-related hyperparameters or priors that govern the influence of the misspecification. We summarise our contributions as follows:
@@
    We introduce RVNP, an amortised SBI method that uses a pre-trained simulator likelihood, an error model, and an importance-weighted autoencoder (burda2015importance) scheme to return robust amortised posterior inference under misspecification without adopting user-tuned misspecification-related hyperparameters or priors. We also introduce RVNP-T, which tunes the final posterior using the simulator and the noise induced by the error model.
@@
    To our knowledge, this is the first example of using amortised variational inference to address the misspecification problem in SBI.
@@ 2.1 SBI formalism
We let $\iota_{\omega}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{l}$, $\bm{x}_{\rm sim}\mapsto\bm{z}_{\rm sim}$ denote a statistical embedding parametrized by $\omega$. This embedding can represent fixed user-defined summary statistics, a pre-defined embedding, or a neural statistic estimator (NSE) where the parameters $\bm{\omega}$ are to be learnt. Lower-dimensional embeddings are important when dealing with high-dimensional data, but can come at the cost of information loss when the embedding is not a sufficient statistic of $\bm{x}_{\rm sim}$ for $\bm{\theta}$ (blum2013comparative). In this paper, we consider examples where we have expert-informed summary statistics, and neural embedding summary statistics.
@@ 2.2 Amortised neural posterior and neural likelihood estimation
The goal of amortised neural posterior estimation is to approximate the unknown posterior distribution $p^{*}(\bm{\theta}|\bm{x_{\rm obs}})$ for all $\bm{x}_{\rm obs}\in\mathcal{X}_{\rm obs}$. After choosing a conditional density estimation architecture $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ parametrized by $\bm{\phi}$, and an architecture for the neural statistic embedding, $\iota_{\omega}$, NPE (2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree) fits for the parameters $\bm{\omega}$ and $\bm{\phi}$ by minimizing the the expected forward Kullback-Leibler (KL) divergence between analytic and approximate posterior
$\mathcal{L}_{\rm NPE}({\bm{\phi},\bm{\omega}})$ $=\mathbb{E}_{p^{*}(x_{\rm obs})}\big[\mathbb{KL}[p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))||q_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big]$ (1)
$=\mathbb{E}_{p^{*}(x_{\rm obs})}\big[\mathbb{E}_{p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))}[-\log q_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big],$
where the expectation is over the unknown true data-generating distribution $p^{*}(\bm{x}_{\rm obs})$. The second line follows from the first because the true unknown posterior and its entropy do not depend on the trainable parameters.
This amortised posterior objective function is not realistic in most situations (schmitt2024detecting), as we rarely have enough real data to approximate the expectation with respect to $p^{*}(\bm{x}_{\rm obs})$, and the true posterior is unknown. Instead, the unknown $p^{*}(\bm{x}_{\rm obs})$ is replaced by the marginal likelihood $p(\bm{x_{\rm obs}})=\int p(\bm{x}_{\rm obs}|\,\bm{\theta})p(\bm{\theta})d\bm{\theta}$. Under this assumption that there is no misspecification, $\bm{x}_{\rm obs}$ can be replaced with $\bm{x}_{\rm sim}$ and the objective becomes
$\mathcal{L}(\bm{\phi},\bm{\omega})_{\rm NPE}:=-\mathbb{E}_{\bm{x}_{\rm sim},\bm{\theta}}[\log q_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm sim}))],$ (2)
where the expectation is over simulator input and output pairs. This is minimised with respect to the parameters $\bm{\omega}$ and $\bm{\phi}$. The success of this objective depends on the assumption that sampling from the evidence is equivalent to sampling from $p^{*}(\bm{x}_{\rm obs})$.
Neural likelihood estimation (NLE) learns a distribution maximising the conditional log-probability of the simulated data
@@
with respect to the flow parameters $\bm{\Psi}$ (where we dropped reference to $\iota_{\bm{\omega}}$), which is equivalent to minimising the KL divergence between the flow and the target distribution. Using neural embeddings for the summary statistics must be implemented in a two-stage manner because it cannot be trained in an end-to-end fashion along with the likelihood proxy (brehmer2020flows). Sampling from the posterior amounts to adopting either a Markov Chain Monte Carlo (MCMC) sampling method or a variational inference algorithm, which will increase the number of architecture choices and hyperparameter tuning, but can recover accurate posterior inference (glockler2022variational).